Find the integral curves of the equation:
dx/(x²+y²) = dy/2xy = dz/{z(x+y)}
1
Expert's answer
2020-03-16T13:25:10-0400

"\frac{dx}{x^2+y^2}=\frac{dy}{2xy}=\frac{dz}{z(x+y)}"


1) "\frac{dx+dy}{x^2+y^2+2xy}=\frac{dz}{z(x+y)}"

"\frac{dx+dy}{(x+y)^2}=\frac{dz}{z(x+y)}, \quad \frac{dx+dy}{x+y}=\frac{dz}{z} ,\quad d(\ln |x+y|)=d(\ln |z|)"

"\ln |x+y|=\ln|z|+C"

"|x+y|=e^C|z|"

"x+y=\hat {C}z"

2) "\frac{dx-dy}{x^2+y^2-2xy}=\frac{dz}{z(x+y)}"

"\frac{dx-dy}{(x-y)^2}=\frac{dz}{ \hat{C} z^2}, \quad d(\frac{-1}{x-y})=d(\frac{-1}{\hat{C}z})"

"\frac{-1}{x-y}=\frac{-1}{\hat{C}z}+C’=\frac{-1}{x+y}+C’"

"\frac{1}{x+y}-\frac{1}{x-y}=\frac{-2y}{x^2-y^2}=C’"


Answer: "x+y=C_1z, \; \frac{-2y}{x^2-y^2}=C_2"